### v1

# PMGS: Reconstruction of Projectile Motion across Large Spatiotemporal Spans via 3D Gaussian Splatting

Abstract
Modeling complex rigid motion across large spatiotemporal spans remains an unresolved challenge in dynamic reconstruction. Existing paradigms are mainly confined to short-term, small-scale deformation and offer limited consideration for physical consistency. This study proposes PMGS, focusing on reconstructing Projectile Motion via 3D Gaussian Splatting. The workflow comprises two stages: 1) Target Modeling: achieving object-centralized reconstruction through dynamic scene decomposition and an improved point density control; 2) Motion Recovery: restoring full motion sequences by learning per-frame SE(3) poses. We introduce an acceleration consistency constraint to bridge Newtonian mechanics and pose estimation, and design a dynamic simulated annealing strategy that adaptively schedules learning rates based on motion states. Futhermore, we devise a Kalman fusion scheme to optimize error accumulation from multi-source observations to mitigate disturbances. Experiments show PMGS’s superior performance in reconstructing high-speed nonlinear rigid motion compared to mainstream dynamic methods.
11footnotetext: These authors contributed equally to this work.22footnotetext: * Corresponding authors: Chu He (chuhe@whu.edu.cn), Dingwen Wang (wangdw@whu.edu.cn) and Lei Yu (ly.wd@whu.edu.cn)
1 Introduction
Dynamic reconstruction has become an engine driving modern film animation, game interaction, and virtual reality. The rise of neural rendering has pushed reconstruction fidelity to unprecedented heights, enabling the depiction of highly challenging deformations, such as subtle tremors in biological tissues Endo . Meanwhile, breakthroughs in generative methods Dream4DGS ; Dreamscene4D have achieved controllable synthesis of dynamic scenes with diverse artistic styles, vastly expanding the boundary of imagination in visual expression.
However, returning to a task related to the essential laws of the physical world: reconstruction of complex rigid motions over large spatiotemporal spans—remains an insufficiently explored challenging issue. The difficulties lie in: 1) Modeling paradigm; 2) Physical consistency.
Most existing modeling paradigms are designed for small range non-rigid motion over short time spans DynaNeRF2 ; DyNeRF ; DeformGS ; 4DGS . These approaches rely on deformation fields with complex temporal regularization to achieve spatiotemporal alignment. However, discrete sampling struggles to capture large nonlinear motion (e.g., accelerated rotation), often resulting in noticeable artifacts and trajectory fragmentation when modeling high-speed rigid motion 4DGS . Additionally, current 4D datasets exhibit similar limitations: in most scenarios, moving objects occupy only a small portion of the scene and exhibit minimal movement. Consequently, many methods prioritize optimizing appearance similarity metrics rather than addressing the core challenge of accurate motion recovery.
Figure 1: Left: Current paradigms focus on small-scale deformation reconstruction. Right: PMGS explores complex rigid motion modeling across large spatiotemporal spans.
The absence of physical consistency poses another major challenge. Many works focus on optimizing appearance similarity while neglecting the pervasive physical constraints in the real world. However, when modeling complex motion across large spatiotemporal spans, relying solely on photometric supervision fails to constrain the vast 3D solution space effectively. For NeRF-based methods NeRFPlayer ; DyNeRF ; DynaNeRF3 , the inherent characteristics of implicit representations encode motion cues within black-box network weights, making physical interactions difficult to model and unintuitive. The introduction of 3D Gaussian Splatting (3DGS) 3DGS has partially addressed this by combining explicit point-based representations with neural rendering, thereby providing a structured carrier for scene understanding. Consequently, some GS-based approaches Endo ; CF ; PEGAUS incorporate off-the-shelf models to enhance physical realism. We argue that this model-stacking strategy introduces compounded uncertainties by layering new approximations over existing ones. More critically, it bypasses the exploration of fundamental physical principles, ultimately reducing the reconstruction task to uninterpretable compensatory computations.
To tackle these issues, we selected a representative task for focused solutions: Projectile motion. As a fundamental concept in mechanics, it describes an object launched with an initial velocity, following parabolic motion under gravity. The study extends from free-fall to complex situations with self-rotation. This phenomenon is not only ubiquitous in real world (e.g., falling objects from heights, throwing events in sports, military ballistic trajectories), while also encompasses the typical challenges of rigid motion, including variable speed, large span, and compound motion (translation and rotation). Consequently, this archetypal paradigm can be generalized to any dynamic scenario within constant force fields in nature, demonstrating universal applicability.
In this paper, we focus on the modeling of complex rigid motions with large spatiotemporal spans, and propose PMGS, a framework that reconstructs Projectile Motion via 3D Gaussian Splatting. By leveraging temporal continuity in video sequences and explicit Gaussian representation, we estimate per-frame SE(3) affine transformation throughout the motion which fundamentally unattainable with NeRF-based methods. Guided by first principles of physics, we introduce an acceleration constraint for motion recovery. Additionally, to alleviate issues of learning oscillations and trajectory fractures caused by fixed training paradigms, we develop a Dynamic Simulated Annealing (DSA) strategy that adaptively schedules the training process based on real-time velocity and displacement variations. Departing from conventional methods that treat physical constraints merely as regularization components, we propose a Kalman filter-based optimal estimation scheme to fuse cross-modal observations thereby minimizing potential disturbances.
Overall, our main contributions can be outlined as:
  • We propose PMGS, a framework integrating target modeling and motion recovery to reconstruct projectile motion. Particularly, we contribute a dataset designed for complex rigid motion modeling at large spatiotemporal scales—an area notably underserved by existing datasets.
  • We introduce an acceleration consistency constraint connecting Newtonian mechanics to pose estimation. Our DSA strategy dynamically adapts training via velocity-displacement motion states, addressing fixed paradigms’ nonlinear motion limitations.
  • We design a Kalman fusion scheme is to dynamically optimize multi-source observations with adaptive weighting, minimizing error accumulation caused by real-world disturbances or training oscillations.
Figure 2: Overview of PMGS. We first segment the target then decompose the motion through centralization to transform the dynamic scene into static. For modeling, we learn a set of Gaussian kernels and align them at the original scale with a set of learnable affine transformation. In the motion recovery, we estimate the target’s SE(3) transformation frame by frame, and comprehensively improve tracking accuracy by integrating physics-enhanced strategies.
2 Related Work
2.1 3D Representation.
Im1 ; Im3 pioneered the exploration of the implicit radiance fields.The milestone work, NeRF NeRF , maps spatial coordinates and viewing directions to color and density through the MLP, enabling differentiable and compact units of complex scenes. However, the implicit representations are inherently constrained in geometric nature, which limits their exploration for editing and interaction. Instead, 3DGS 3DGS introduces an shift from implicit to explicit paradigm, the point-based representation greatly expands its potential of scene understanding.
2.2 Dynamic Neural Rendering.
Current dynamic neural rendering include: (1) Point deformation fields DynaNeRF2 map sampled points at each timestep to a static canonical space. (2) Time-aware Volume Rendering DyNeRF ; NeRFPlayer employs factorized voxels to independently compute features for each point. (3) The Gaussian deformation field DeformGS ; MotionGS ; S4d transforms 3D Gaussians to their target position at specific timesteps. The critical issues lie in the deformation field struggles to model large inter-frame displacements that commonly occur in real-world scenarios, which can be further validated by experiments on Panoptic Studio dataset Sportsdatasest conducted by 4DGS 4DGS . Moreover, due to the lack of physical consistency constraints, the network tends to focus on fitting appearance similarity while neglecting structural accuracy at the geometric level. As a typical example, when there are large areas of nearly uniform color, Gaussian bodies will float chaotically within regions of similar color Dynamic3DGS .
2.3 Enhancing GS with Physical Assistance.
Introducing physical assistance to enhance 3DGS has been widely applied. CF ; Endo ; FSGS utilize off-the-shelf monocular depth estimation models DPT for scene initialization or sparse reconstruction. Spring devised a Spring-Mass model to simulate the elastic objects’s falling and collision. PEGAUS is designed for 3D dataset generation which integrates the PyBullet engine to emulate the placement of objects and their dynamic processes. Dreamscene4D ; SAGS enhances pose estimation by ensuring scale consistency and local rigidity. Essentially, current approaches use rough approximations from external physics models to fix ill-posed inverse problems in vision systems. Nevertheless, pre-trained engines lose confidence sharply in unusual scenarios, and their black-box makes errors untraceable—causing mistakes to snowball. Differently, we propose to explore strategies grounded directly in first principles of physics particlegs , rather than relying on external models for unreliable assistance.
3 Method
PMGS takes a monocular video as input to reconstruct the target and achieve full-sequence projectile motion recovery. For modeling appearance and geometry, we equivalently converts dynamic scenes into static through motion decomposition, and combined with the improved point density control strategy to enhance the geometric accuracy. During motion recovery, we leverage video temporal continuity and the explicit Gaussian representation to estimate per-frame SE(3) transformations. To ensure physically consistent pose learning, we introduce an acceleration consistency constraint that establishes direct connections between pose estimation and Newtonian dynamics priors. Furthermore, we devise a DSA strategy, which adaptively schedules the optimization process according to real-time motion states. Finally, we design a Kalman fusion strategy to optimize the error accumulation across multi-modal observation sources.
3.1 Target modeling
To model the target’s appearance and geometry, we reformulate the dynamic scene as a static object-centralized scene.
Centralization. We firstly use the pre-trained SAM SAM to separate the background and the dynamic target. Then, by establishing an object-centralized normalized coordinate Dreamscene4D , we decompose the complex projectile motion: the object’s autorotation is disentangled into pseudo multi-view observations, while translational displacements are eliminated.
Reconstruction. Following the 3DGS 3DGS , we use a set of Gaussian kernels to explicitly reconstruct the appearance and geometry of the object.
Notably, geometric quality directly affects motion recovery. Therefore, the structural flaws of Gaussian representation cannot be ignored: (1) Oversize Gaussians. They frequently occur at boundaries to incorrectly fit the viewpoint variations, inducing jagged edges and spiky artifacts. (2) Discretized Gaussians. They appear distant from the target, distorting centroid computation and introducing systematic errors in motion estimation. To address these, we propose an improved point density control strategy: (1) Axial constraint: Hard pruning of Gaussians with excessive axial lengths; (2) Outlier removal: Gaussians whose distance to their nearest neighbors exceeds the average value are removed. This can be mathematically expressed as:
$\mathcal{G}_{\text{pruned}}=\left\{g_{i}\in\mathcal{G}\,\middle|\,\begin{aligned} &\max\big{(}\mathrm{eig}(\Sigma_{i})\big{)}\leq\tau_{L}\quad\cap\\ &\min_{g_{j}\in\mathcal{G}}\|\mu_{i}-\mu_{j}\|_{2}\leq\tau_{D}\cdot D_{avg}\end{aligned}\right\}$ (1)
where $\Sigma_{i}$ and $\mu_{i}$ represent the covariance matrix and position of a Gaussian $g_{i}$ respectively. For axial constraint, Gaussians with length of principal axis $\max\big{(}\mathrm{eig}(\Sigma_{i})\big{)}>\tau_{L}$ are discarded, where $\tau_{L}$ is a predefined threshold controlling shape anisotropy. Besides, an outlier is defined as a Gaussian whose minimum distance to other is much larger than the average distance $D_{avg}=\frac{1}{|\mathcal{G}|^{2}}\sum_{m,n}\|\mu_{m}-\mu_{n}\|_{2}$, where $\tau_{D}$ is a distance-based filtering factor.
During the optimization, the loss $\mathcal{L}_{GS}$ is used to quantify the difference between the real image $\hat{I}$ and render image $I$:
$\mathcal{L}_{GS}=(1-\lambda)\lVert{\hat{I}-I}\rVert_{1}+\lambda\mathcal{L}_{D-SSIM}$ (2)
Up to this point, we consequently obtain a Gaussian field $G_{0}$ at the centralized scale.
Registration. Finally, we register the static Gaussian $G_{0}$ at the centralized scale to the original dynamic scene, thus facilitating subsequent motion recovery. This can be achieved via a set of learnable affine transformations $T_{reg}=[\mathbf{r,t},s]$, i.e., quaternion rotation $\mathbf{r}\in\mathbb{R}^{4}$, translation vector $\mathbf{t}\in\mathbb{R}^{3}$, and constant scalar scale $s$. Then the aligned Gaussian field $G_{1}$ can be represented as $T_{reg}\odot G_{0}$. In addition to $\mathcal{L}_{GS}$, we further apply VGG loss $\mathcal{L}_{VGG}$ and grid-based total variation loss $\mathcal{L}_{TV}$ Spring ; 4DGS to achieve a better awareness for accurate registration. The total loss can be formulated as:
$\mathcal{L}_{Align}=\lambda_{1}\mathcal{L}_{GS}+\lambda_{2}\mathcal{L}_{VGG}+\lambda_{3}\mathcal{L}_{TV}$ (3)
The optimization target is to minimize $\mathcal{L}_{Align}$ between the rendered image of $G_{1}$ and the first frame $\hat{I}_{0}$ of original dynamic scene:
$T_{reg}=\mathrm{arg}\underset{T_{reg}}{\mathrm{min}}\mathcal{L}_{Align}(\mathcal{R}(T_{reg}\odot G_{0}),\hat{I}_{0})$ (4)
Finally, we obtain this set of Gaussian $G_{1}$, which can well-represent the appearance and geometry of the object.
Figure 3: DSA strategy. Right: An object accelerates in a constant gravity field, with different velocity $v$ and displacements $s$ corresponding to each timestamp. Left: Red curve represents the initial learning rate. Blue points shows the final learning rate after exponential decay.
3.2 Physics-Enhanced Motion recovery
During the dynamic phase, we estimate 6DoF pose changes of the object in total $(n+1)$ frames, which corresponds to a sequence of temporal SE-3 transformations $T_{n}=[\mathbf{r}_{n},\mathbf{t}_{n}]$. We introduce an acceleration consistency constraint to establish a direct connection between pose learning and physical priors. Additionally, we incorporate optical flow smoothing and a DSA training strategy to improve training stability.
Acceleration consistency constraint. The object moves in a constant force field, therefore its acceleration remains constant. Based on this, we first calculate the object’s center of mass $\sigma$. Since the Gaussian have been processed isotropically, the spatial coordinates of $\sigma$ can be computed as:
$\sigma=\frac{\sum_{i=1}^{N}R_{i}^{3}\mu_{i}}{\sum_{i=1}^{N}R_{i}^{3}}$ (5)
where $N$ represents the total number of Gaussian kernels, $R$ is the radius of each kernel, and $\mu_{i}=[x_{i},y_{i},z_{i}]$ denotes the spatial position of the kernel. In turn, the acceleration of the object at the current moment $t$ can be calculated as:
${a}^{t}=\frac{[\sigma^{(t+\Delta{t})}-\sigma^{t}]-[\sigma^{t}-\sigma^{(t-\Delta{t})}]}{(\Delta t)^{2}}$ (6)
where $\Delta{t}$ denotes the time interval between two adjacent frames. Subsequently, the acceleration $\textbf{a}^{t}$ can be decomposed into components parallel and orthogonal to the gravity direction g:
$\mathbf{a}^{t}_{\parallel}$ $=(\mathbf{a}^{t}\cdot\mathbf{g})\,\mathbf{g},$ (7)
$\mathbf{a}^{t}_{\perp}$ $=\mathbf{a}^{t}-(\mathbf{a}^{t}\cdot\mathbf{g})\,\mathbf{g}$
According to the Newtonian dynamics, there should be a constant acceleration along the gravity direction, while the components orthogonal to it should remain zero:
$\mathcal{L}_{Acc}=\lVert{\mathbf{a}^{t}_{\parallel}+(a_{\perp}^{(t+\Delta{t})}-a_{\perp}^{t})}\rVert_{2}^{2}$ (8)
Building upon Eq.8, we are able to effectively regularize the learning of translation components across consecutive frames, even under scale ambiguity, thereby ensuring the physical consistency of motion recovery without requiring absolute values such as gravitational acceleration.
Dynamic Simulated Annealing. As shown in Figure 3, a fixed $lr_{init}$ struggles to adapt to varying object’s velocities: excessively large $lr_{init}$ induce training oscillations during low-speed phases, whereas small $lr_{init}$ hinder convergence when tracking high-speed targets. To address this velocity-displacement coupling effect, we dynamically schedule the $lr_{init}$ according to real-time displacement.
Specifically, for a moving object with an initial velocity $v_{0}$ and constant acceleration $a$, the displacement during the interval from timestamp $t$ to $(t+\Delta t)$ can be derived as:
$\Delta s_{t}=(a\Delta t)t+[\frac{1}{2}a(\Delta t)^{2}+v_{0}\Delta t]$ (9)
As the initial velocity $v_{0}$, acceleration $a$ and the time interval $\Delta t$ are constant value, thus Equation 9 establishes the linear correlation between object’s displacement $\Delta s$ and timestamp $t$, i.e., $\Delta s\propto t$. This temporal-displacement relationship therefore requires proportional scaling of the $lr_{init}$ with displacement. Therefore, we implement scheduling where the learned displacement from the preceding timestep governs the subsequent $lr_{init}$, as visualized in Figure 3.
As the iteration progresses, the Gaussians gradually approach the target spatial position. Correspondingly, the learning rate needs to be reduced so that the step gradually shrinks for finetuning. We utilized the commonly used exponential decay to achieve this. In addition, for timestamps with large displacements, we appropriately extend the number of iterations to better find the optimal solution.
In summary, we follow the DSA strategy for frame-by-frame tracking of the pose $T_{n}$ during motion recovery, and the total optimization is to minimize a composite of losses:
$T_{n}$ $=\mathrm{arg}\underset{T_{n}}{\mathrm{min}}(\lambda_{1}\mathcal{L}_{GS}(\mathcal{R}(G_{n}),\hat{I}_{n})+\lambda_{2}\mathcal{L}_{Acc})$ (10)
3.3 Cross-modal Kalman fusion
Furthermore, we introduce optical flow smoothing computed based on Optical for emphasizing attention to motion variations in long-term tracking. Though some works MotionAware also incorporate optical flow or priors as auxiliaries, but they fail to account for the inaccurate optical flow estimation caused by degraded textures under motion blur, which in turn introduces negative effects. Differently, we design a data fusion strategy based on Kalman Filter Kalman to dynamically balance the weights of observation sources, suppress cumulative error, and update the optimal estimate in real-time.
Step1-Definition: The basic elements in Kalman fusion namely the system prediction model and observations. When applied to our motion recovery process, these components correspond to the following:
  • System model: Displacement prediction based on the acceleration consistency constraint (Eq.8).
  • Observation 1: The displacement calculated via back-projection from the inter-frame optical flow field $z_{flow}^{t}$.
  • Observation 2: The displacement of the current frame actually learned by the network $z_{learn}^{t}$.
Step2-Prediction: In the prediction step, we use the system dynamic model and previous state estimation to predict the current state, which can be formulated as:
$\textbf{X}_{t|t-1}=\begin{bmatrix}\Delta s_{t|t-1}\\ v_{t|t-1}\end{bmatrix}=\textbf{F}\textbf{X}_{t-1|t-1}+\textbf{B}a_{k}+\textbf{w}_{k},$ (11)
$\textbf{F}=\begin{bmatrix}1&\Delta t\\ 0&1\end{bmatrix},\textbf{B}=\begin{bmatrix}\frac{1}{2}(\Delta t)^{2}\\ \Delta t\end{bmatrix},\textbf{Q}=\begin{bmatrix}\sigma_{\Delta s}^{2}&0\\ 0&\sigma_{v}^{2}\end{bmatrix}$
where $\textbf{X}_{t|t-1}$ is the prediction based on the optimal estimate at previous timestamp, F is the state transition matrix, B is the control input matrix, and $a_{t}$ is the control input calculated from Eq.6, which approximates to $a_{t-1}$ as the system model defined. $\textbf{Q}=\mathbb{E}[\textbf{w}_{k}\textbf{w}_{k}^{T}]$ is the covariance matrix of the process noise $\textbf{w}_{k}$ (includes noise $\sigma_{\Delta s}^{2}$ and $\sigma_{v}^{2}$) that describes the uncertainty of the model. The prediction covariance then can be calculated as:
$\textbf{P}_{t|t-1}=\textbf{F}\textbf{P}_{t-1|t-1}\textbf{F}^{T}+\textbf{Q}$ (12)
where $\textbf{P}_{t-1|t-1}$ is the posterior covariance of the previous state.
[…]
$\textbf{z}_{t}=\begin{bmatrix}z_{flow}^{t}\\ z_{learn}^{t}\end{bmatrix}=\textbf{H}\textbf{X}_{t|t}+\textbf{v}_{k},$ (13)
$\textbf{H}=\begin{bmatrix}1&0\\ 1&0\end{bmatrix},\textbf{R}=\begin{bmatrix}\sigma_{flow}^{2}&0\\ 0&\sigma_{learn}^{2}\end{bmatrix}$
where H is the observation matrix, $\textbf{R}=\mathbb{E}[\textbf{v}_{k}\textbf{v}_{k}^{T}]$ is the covariance of noise $\textbf{v}_{k}$.
Step4-Update: Based on above content, we update the Kalman gain $\textbf{K}_{t}$ and the posterior covariance $\textbf{P}_{t|t}$:
$\textbf{K}_{t}=\textbf{P}_{t|t-1}\textbf{H}^{T}(\textbf{H}\textbf{P}_{t|t-1}\textbf{H}^{T}+\textbf{R})^{-1}$ (14)
$\textbf{P}_{t|t}=(\textbf{I}-\textbf{K}_{t}\textbf{H})\textbf{P}_{t|t-1}$ (15)
Finally, we can obtain the optimal estimated value at the current timestamp:
$\textbf{X}_{t|t}=\textbf{X}_{t|t-1}+\textbf{K}_{t}(\textbf{z}_{t}-\textbf{H}\textbf{X}_{t|t-1})$ (16)
Throughout the fusion process, we fully integrate the physical model, optical flow observations, and network outputs. The Kalman gain is used to automatically adjust the weights, thereby minimizing the error accumulation across multi-modal observation sources. Therefore, even when real-world disturbances exist or poor image quality compromises texture observations Motionblurgs ; comogs , this strategy maintains strong robustness.
4 Experiments
Figure 4: Qualitative comparison on both synthetic and real datasets. PMGS generalizes well across different scenarios and accurately reconstructs full-sequence motion. The displayed results are uniformly sampled, for complete and coherent motion recovery, please refer to the video on the project page.
Datasets Source Type Phy. Mono.
LaSOT lasot Real Rigid $\times$ $\times$
D-NeRF DynaNeRF2 Syn. Def. $\times$ $\checkmark$
HyperNeRF DynaNeRF3 Real Def. $\times$ $\checkmark$
Neural3D DyNeRF Syn. Def. $\times$ $\times$
PEGASUS PEGAUS Syn. Rigid PyBullet $\checkmark$
SpringGS Spring 6S + 3R Def. Hooke’s Law $\times$
PMGS 6S + 3R Rigid Newton’s Law $\checkmark$
Table 1: Datasets comparison (Phy.-Whether physical prior is introduced; Mono.-Whether monocular; Syn.-Synthetic; Def.-Deformable).
Grenade Missile Cola Doll Chocolate Rugby Box Shark Bear Mean
[…]
2DGS 24.82 0.895 0.153 17.42 0.714 0.325 15.97 0.847 0.195 17.99 0.792 0.218 15.19 0.758 0.275 21.56 0.830 0.185 19.72 0.973 0.172 11.57 0.794 0.101 11.78 0.823 0.201 17.34 0.825 0.202
InstantSplat 25.44 0.900 0.139 17.58 0.724 0.300 17.03 0.852 0.175 18.10 0.797 0.221 16.10 0.761 0.255 24.21 0.877 0.165 20.48 0.977 0.152 12.88 0.810 0.097 11.79 0.823 0.090 18.18 0.835 0.177
CFGS* 21.82 0.873 0.231 13.70 0.675 0.259 14.47 0.845 0.284 18.00 0.788 0.307 11.93 0.739 0.317 21.37 0.822 0.196 17.19 0.946 0.162 11.32 0.742 0.271 11.54 0.772 0.212 15.70 0.800 0.248
4DGS 20.52 0.907 0.112 22.60 0.829 0.082 20.25 0.859 0.154 23.92 0.947 0.063 17.91 0.934 0.182 Failed Failed Failed 20.19 0.874 0.131 15.32 0.709 0.302 Failed Failed Failed 19.11 0.866 0.156
MotionGS 22.64 0.876 0.166 19.14 0.857 0.156 15.39 0.823 0.217 22.70 0.854 0.119 16.99 0.902 0.115 17.37 0.855 0.151 15.92 0.512 0.464 10.38 0.654 0.353 13.99 0.887 0.231 17.59 0.782 0.227
DynamicGS 36.32 0.971 0.064 26.37 0.906 0.032 28.44 0.939 0.026 31.81 0.942 0.050 25.36 0.897 0.018 Failed Failed Failed 18.97 0.704 0.152 24.92 0.952 0.085 Failed Failed Failed 28.03 0.901 0.058
PMGS 33.58 0.967 0.011 28.87 0.915 0.025 26.50 0.940 0.014 34.38 0.959 0.014 25.01 0.924 0.020 33.72 0.954 0.012 30.45 0.974 0.019 25.07 0.952 0.094 26.64 0.959 0.059 29.35 0.949 0.029
Table 2: Comparison of video reconstruction.
Grenade Missile Cola Doll Chocolate Rugby Box Shark Bear Mean
[…]
CFGS* 0.249 0.550 0.602 0.203 0.129 0.218 0.154 0.533 0.563 0.310 0.358 0.459 0.318 0.574 0.610 0.094 0.562 0.604 0.079 0.392 0.425 0.252 0.425 0.458 0.199 0.319 0.364 0.206 0.426 0.478
4DGS 0.359 0.661 0.703 0.608 0.194 0.248 0.354 0.630 0.661 0.401 0.563 0.621 0.287 0.668 0.687 Failed Failed Failed 0.144 0.553 0.592 0.141 0.505 0.556 Failed Failed Failed 0.302 0.549 0.589
MotionGS 0.288 0.474 0.505 0.541 0.655 0.704 0.423 0.511 0.551 0.501 0.505 0.559 0.549 0.501 0.522 0.249 0.643 0.668 0.299 0.267 0.313 0.298 0.557 0.629 0.179 0.526 0.573 0.414 0.495 0.540
DynamicGS 0.548 0.227 0.269 0.910 0.118 0.156 0.898 0.182 0.226 0.663 0.311 0.363 0.995 0.070 0.090 Failed Failed Failed 0.262 0.439 0.470 0.679 0.094 0.155 Failed Failed Failed 0.707 0.205 0.247
PMGS 0.998 0.168 0.254 0.995 0.091 0.130 0.997 0.069 0.080 0.993 0.092 0.152 0.987 0.168 0.214 0.952 0.021 0.148 0.940 0.055 0.086 0.876 0.021 0.093 0.969 0.043 0.047 0.967 0.080 0.133
Table 3: Comparison of motion recovery.
4.1 Experimental Settings
Datasets. We constructed both synthetic and real-world dataset. For synthetic data, we collected 6 models from public 3D communities Blender , then set up a constant gravity field in Blender. The objects were launched with a randomly set initial velocity and accompanied by autorotation. A fixed camera was used to capture the dynamic scene at 120 FPS, rendering 120 images at a resolution of 1024x1024. For the real-world dataset, we threw 3 different objects and captured the scene using a fixed camera, with an exposure time of 1/2500s, a frame rate of 60 FPS, and a resolution of 4128×2752. During training, the images were resized to below 1600 pixels. Due to the frame rate limitations of the imaging device and the real-world projectile objects do not undergo sufficiently complete rotations (i.e., failing to exhibit a full 360° surface), we captured three projectile sequences for each model from the same viewpoint. Samples were then randomly selected from these sequences to perform modeling.
For reconstruction, the training/validation split strictly adheres to 3DGS. As for the motion recovery, we focus on recovering the full-sequence motion (all frames), thus no split is required. All competitors follow the aforementioned dataset division for fair metric calculation.
Furthermore, we conducted a comprehensive comparison between our proposed dataset and existing 4D datasets, as detailed in Table 1. PMGS covers both real and synthetic sources and sets the challenging monocular scenario, focusing on long-duration, large-span complex rigid motions. All motions strictly adhere to real-world physical laws, which has not been considered in many synthetic datasets.
Metrics. We evaluate video reconstruction and motion recovery separately. For video reconstruction, we use PSNR, SSIM, and LPIPS SSIM ; PSNR . For motion recovery, we compute the bounding boxes of the target object in both the real and rendered image, and then calculate the IoU, absolute trajectory error (ATE) and RMSE IoU ; Tracking ; CF to comprehensively assess the spatial accuracy of tracking and trajectory. It should be clarified that, all evaluations were computed after fully removing backgrounds.
Implementation Details. We conducted experiments in the PyTorch framework with an Nvidia RTX 4090 GPU. During the centralization phase, we cropped the target from the original data and placed it at the center of a 512x512 canvas. In the registration phase, we trained a set of $[\mathbf{r,t},s]$ to align the static Gaussian with the first frame of the dynamic scene within 10,000 iterations, using a learning rate of 5e-5. This process is fast, typically taking only 2 minutes. In the motion recovery phase, the initial base learning rate applied to the frame with the minimal velocity was set to 1.0e-3, with a base iteration count of 1000. The weights for the $\mathcal{L}_{GS}$, $\mathcal{L}_{Acc}$ and $\mathcal{L}_{Smooth}$ were set to 0.7, 0.2 and 0.1.
4.2 Comparison
We select 2DGS 2DGS , InstantSplat Instantsplat as static competitors. Dynamic competitors include 4DGS 4DGS , DynamicGS Dynamic3DGS and MotionGS MotionGS . Furthermore, we employed CFGS CF as a baseline by disabling its depth estimation for reconstruction, and instead inputting our pre-trained Gaussian model for purely 6DoF pose estimation comparison (symbolized as CFGS*).
Comparison of video reconstruction. We report the quantitative results in Table 2. Our algorithm demonstrates robust performance, as the proposed density control strategy effectively enhances appearance and geometric quality, thereby providing a solid foundation for dynamic reconstruction. CFGS underperforms, primarily due to inherent flaws in depth inference pipelines. For object-specific modeling under non-open scenarios, the monocular observation system yields an underdetermined solution manifold, where maintaining 3D consistency over extended spatiotemporal intervals becomes theoretically unattainable.
Comparison of motion recovery. We uniformly sample frames from complete sequences and visualize in Figure 4, and the quantitative results are reported in Table 3. PMGS performs well in all three tracking metrics, demonstrating accurate estimation of both translation and rotation. 4DGS and MotionGS exhibit obvious trajectory fractures and artifacts—typical limitations of dynamic algorithms when modeling large-scale motions. While DynamicGS demonstrates promising performance, it exhibits limited generalization capability to certain instances. For CFGS, even with our well-trained 3D model provided, the spatial position of Gaussian is incorrectly driven toward infinity under only photometric supervision to forcibly fit image similarity. These findings collectively reveal the effectiveness of physics constraints and adaptive training scheduling for recovering complex motions across large spatiotemporal domains.
Additional experimental results and extended applications are presented in Figure 5, showing the potential of PMGS in understanding and interacting with the real world.
Figure 5: Extended experiments. (a) Novel views synthesis. (b) Multi-objects interaction. (c) Future motion trajectory prediction based on the 3D centroid coordinates with MLP. (d) Instantaneous displacement percpetion.
Video reconstruction Motion recovery
[…]
CFGS* 16.82 0.799 0.249 0.206 0.426 0.478 103 mins
Model 1 (w/o density control) 24.61 0.910 0.089 0.635 0.334 0.791 51 mins
Model 2 (w/o $L_{Acc}$ DSA) 25.63 0.929 0.045 0.886 0.137 0.188 64 mins
Model 3 (w/o Kalman) 28.95 0.940 0.035 0.933 0.098 0.147 42 mins
PMGS (Ours) 29.35 0.949 0.029 0.967 0.080 0.133 48 mins
Table 4: Ablation study of different models.
Figure 6: Ablation results for different comparison models. PMGS in its complete form exhibits superior stability.
4.3 Ablation Study
We conducted ablation experiments to verify the effectiveness of different novel modules.
Effectiveness of point density control. As evidenced in Table 4, Model 1 exhibits degradation, demonstrating that poor reconstruction directly compromises tracking accuracy. Incorrect geometry and appearance representations may cause target mislocalization—especially when floating Gaussian points distort the calculation of central points and bounding boxes, thus resulting in cumulative errors over time.
Effectiveness of physical constraints. Model 2 ablates $L_{Acc}$ and DSA strategy, retaining only photometric supervision and fixed training paradigm. As shown in Figure 6, the spatial position of the object exhibits drift, and errors are observed obviously in rotational orientation. The removal of DSA significantly increases the optimization time cost, revealing the necessity of adaptive adjustment of the learning rate under high-speed displacement.
Effectiveness of Kalman fusion. The removal reduced computational cost overhead at the expense of a decrease in stability. In more challenging scenarios (e.g., under substantial time-varying force interference), the strategy’s role in ensuring robustness would be more pronounced.
5 Conclusion
In this paper, we propose PMGS for reconstructing projectile motion over large spatiotemporal scales from monocular videos. Through a two-stage pipeline involving dynamic scene decomposition modeling and physically enhanced motion recovery, we effectively alleviate the issues of trajectory fragmentation and physical implausibility encountered by existing dynamic neural rendering methods when handling high-speed, non-linear rigid motion. Moreover, the proposed cross-modal Kalman fusion optimizes error accumulation from multi-source observations, further enhancing robustness against real-world disturbances. PMGS provides a novel perspective for the explicit modeling and understanding of the 3D physical world. In the future, we plan to delve into more challenging topics, such as achieving accurate 3D motion reconstruction under low-quality imaging conditions or changing force field environments.